geerky42 (geerky42):

a(x+7) + b(x+5) = x+1 Find a - b

geerky42 (geerky42):

\[\Huge x(a+b) + 7a + 5b = x+1\]Now what?

OpenStudy (sirm3d):

the constants are equal; the coefficients of x are equal, that is \[\huge 7a+7b=1\] \[\huge a+b=1\]

OpenStudy (sirm3d):

correction: the first equation is \[\huge 7a+5b=1\]

geerky42 (geerky42):

And a+b = 1 too?

OpenStudy (sirm3d):

yes, that's right.

geerky42 (geerky42):

So a - b = -5?

OpenStudy (sirm3d):

yup.

geerky42 (geerky42):

Well, that was easier than I excepted, but what if I have something like this: x(a+b)+7a+5b= -1(x+1)

geerky42 (geerky42):

-(a+b) = 1 -(7a + 5b) = 1 Right?

geerky42 (geerky42):

@sirm3d

OpenStudy (sirm3d):

right. just compare coefficients of x and constants.